# Efficient Learning for Linear Properties of Bounded-Gate Quantum Circuits

### v1

@@ A.4 Literature review
$\rho_{\Lambda}=\sum_{\bm{\omega}\in\mathfrak{C}(\Lambda)}\Phi_{\bm{\omega}}(\bm{x})\rho_{\bm{\omega}}\leavevmode\nobreak\ \textnormal{and}\leavevmode\nobreak\ \rho=\sum_{\bm{\omega}\in\mathfrak{C}(d)}\Phi_{\bm{\omega}}(\bm{x})\rho_{\bm{\omega}},$ (85)
@@
$\Psi_{\bm{\omega}_{i}}(\bm{x}_{i})\equiv\nabla_{\bm{x}_{i}}\Phi_{\bm{\omega}_{i}}(\bm{x}_{i})=\begin{cases}0\leavevmode\nobreak\ &\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=0\\ -\sin(\bm{x}_{i})&\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=1\\ \cos(\bm{x}_{i})&\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=-1\end{cases}.$ (89)
@@
where Eqs. (90h)-(90h) are obtained by taking the expectation value on the gradient term $\Psi_{\bm{\omega}_{i}}$ for $\forall i\in[d]$, the quantity $\alpha_{\pm 1,\bm{\omega}_{2:d}}$ refers to $\operatorname{Tr}(\rho_{\pm 1,\bm{\omega}_{2:d}}O)$ with the frequency value at the $1$-st position being ±$\pm 1$ (the same rule applies to other terms such as $\alpha_{\bm{\omega}_{1:i},i+1,\bm{\omega}_{i+2:d}}$), and Eq. (90j) is obtained by taking integral over each term in the bracket and the orthogonality of basis functions in $\Phi_{\bm{\omega}}(\bm{x})$ and $\Phi_{\bm{\omega}^{\prime}}(\bm{x})$ as shown in Eq. (87).
@@
Note that depending on the number of non-zero entries of the frequency (i.e., $\|\bm{\omega}\|_{0}$), the left and right-hand sides in Eq. (92) can be decomposed into $d-\Lambda$ parts and $d-1$ parts, respectively. Mathematically, the left hand side of Eq. (92) yields
@@
$\bigstar$=★ = $\Biggl{[}\bigg{(}\sum_{\bm{\omega}_{2:d},\|\bm{\omega}_{2:d}\|_{0}=0}2^{-\|[1,\bm{\omega}_{2:d}]\|_{0}}|\alpha_{1,\bm{\omega}_{2:d}}|^{2}+\cdots+\sum_{\bm{\omega}_{2:d},\|\bm{\omega}_{2:d}\|=d-1}2^{-\|[1,\bm{\omega}_{2:d}]\|_{0}}|\alpha_{1,\bm{\omega}_{2:d}}|^{2}\bigg{)}$ (96)
+$\cdots$+ ⋯
@@
$\Lambda^{\prime}\sum_{\bm{\omega},\|\bm{\omega}\|_{0}=\Lambda^{\prime}}2^{-\|\bm{\omega}\|_{0}}|\alpha_{\bm{\omega}}|^{2},\quad\text{with}\leavevmode\nobreak\ \Lambda<\Lambda^{\prime}<d.$ (101)
@@
Figure F.5: A visual interpretation about the appearance times of $\tilde{\bm{\omega}}$ in the right-hand side of Eq. (100).
Combining with the above two equations, we know that the total number of frequencies in Eq. (101) and the right-hand side of Eq. (100) is the same. Accordingly, their equivalence can be reached if we can show that each frequency $\bm{\omega}\in\{0,\pm 1\}^{d}$ with $\|\bm{\omega}\|_{0}=\Lambda^{\prime}$ appears $\Lambda^{\prime}$ times in the right-hand side of Eq. (100). This is indeed the case, as the visual interpretation is shown in Fig. F.5. Denote that the specified frequency as $\tilde{\bm{\omega}}$ with $\|\tilde{\bm{\omega}}\|_{0}=\Lambda^{\prime}$. In other words, there are $d-\Lambda^{\prime}$ entries whose value is 00. This property allows us to determine how many times $\tilde{\bm{\omega}}$ appears on the right side of Eq. (100). In particular, as shown in the left panel of Fig. F.5, the right-hand side of Eq. (100) can be divided into $d$ groups, depending on the location of ±$\pm 1$. Note that among these $d$ groups, every entry with the value 00 in $\tilde{\bm{\omega}}$ precludes one group, and only $d-(d-\Lambda^{\prime})=\Lambda^{\prime}$ feasible groups are preserved, as shown in the right panel of Fig. F.5. Moreover, for each feasible group, $\tilde{\bm{\omega}}$ can only appear once, because of the orthogonality of different frequencies. Consequently, we obtain that each frequency $\tilde{\bm{\omega}}\in\{0,\pm 1\}^{d}$ with $\|\tilde{\bm{\omega}}\|_{0}=\Lambda^{\prime}$ appears $\Lambda^{\prime}$ times, indicating that the right-hand side in Eq. (100) equals to Eq.  (101).
@@
where $s_{j}^{(\bm{x}^{(i)})}$ denotes the randomized measurement outcome for the $j$-th qubit for the state $\rho(\bm{x}^{(i)})$ with $s_{j}^{(\bm{x}^{(i)})}\in\{\ket{0},\ket{1},\ket{\pm},\ket{\pm\imath}\}$, Eq. (108b) uses the fact that each $\bm{x}^{(i)}$ is sampled independently and uniformly from $[-\pi,\pi]^{d}$, Eq. (108c) employs the explicit formula of the truncated trigonometric monomial kernel $\kappa_{\Lambda}$, Eq. (108f) adopts trigonometric expansion of the quantum state, i.e., $\rho(\bm{x}^{(i)})=\sum_{\bm{\omega}}\Phi_{\bm{\omega}}(\bm{x}^{(i)})\rho_{\bm{\omega}}$, and Eq. (108h) comes from the orthogonality of basis functions in Eq. (87).
@@
where the first inequality uses the Cauchy–Schwarz inequality and $|\Phi_{\bm{\omega}}(\bm{x})|\leq 1$, the second inequality adopts the von Neumann’s trace inequality with Hölder’s inequality, and the last inequality exploits $\|\tilde{\rho}_{1}(\bm{x})\|_{1}\leq 3^{K}$ [82, Eq. (F37)] and the condition $\|O\|_{\infty}\leq B$.
@@
Denote the set of truncated frequencies as $\mathfrak{C}(\Lambda)=\{\bm{\omega}|\bm{\omega}\in\{0,\pm 1\}^{d},\leavevmode\nobreak\ s.t.\leavevmode\nobreak\ \|\bm{\omega}\|_{0}\leq\Lambda\}$. This leads to
@@
The proposed ML model, as mentioned in the main text, is not only sample efficient but also computationally efficient when the quantity $C$ is well bounded.Here we analyze the computational cost of the proposed ML model. Recall that our proposal consists of two stages: training dataset collection and model inference. As a result, to demonstrate the efficiency of our proposal, it is equivalent to exhibiting the computational efficiency at each stage. In the following, we separately analyze the computational time of our proposal required in the training and inference.
@@ H.1 Implementation of the classical prediction model
Figure H.6: Purely classical learning model with the full expansion. The pipeline is similar to the one presented in Fig. 1 of the main text. The only difference is the way of collecting training data, where the circuit layout is known by the learner and the expectation value $\operatorname{Tr}(\rho(bx)O)$ should be efficiently calculated by the employed classical simulators.
The implementation of the protocol introduced in Corollary 1 is visualized in Fig. H.6. The only difference with the one introduced in the main text is the way of collecting training data. Specifically, when the tunable quantum circuit architecture formed by $\operatorname{RZ}+\operatorname{\text{CI}}$ is known to the learner, the classical simulators (e.g., tensor network simulators, LOWESA, and near Clifford circuits simulators) are employed to collect the training data
@@
$h_{\mathsf{c}}(\bm{x},O)=\frac{1}{n}\sum_{i=1}^{n}\kappa\left(\bm{x},\bm{x}^{(i)}\right)\operatorname{Tr}(\rho(\bm{x}^{(i)})O)\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ \kappa(\bm{x},\bm{x}^{(i)})=\sum_{\bm{\omega},\|\bm{\omega}\|_{0}\leq d}2^{\|\bm{\omega}\|_{0}}\Phi_{\bm{\omega}}(\bm{x})\Phi_{\bm{\omega}}(\bm{x}^{(i)})\in\mathbb{R}.$ (125)
@@ H.2 Proof of Corollary 1
[…]
@@
$\rho(\bm{x})=U(\bm{x})(\ket{0}\bra{0})^{\otimes N}U(\bm{x})^{\dagger}=\sum_{\bm{\omega}}\Phi_{\bm{\omega}}(\bm{x})\langle\langle 0|\operatorname{\mathfrak{U}}_{\bm{\omega}}^{\dagger}\equiv\sum_{\bm{\omega}}\Phi_{\bm{\omega}}(\bm{x})\rho_{\bm{\omega}},$ (137)
where the trigonometric monomial basis is $\Phi_{\bm{\omega}}(\bm{x})=\prod_{i=1}^{d}\begin{cases}1\leavevmode\nobreak\ &\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=0\\ \cos(\bm{x}_{i})&\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=1\\ \sin(\bm{x}_{i})&\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=-1\end{cases}$, then we can form the state prediction model in Eq. (75), i.e.,
$\hat{\sigma}_{n}(\bm{x})=\frac{1}{n}\sum_{i=1}^{n}\kappa\left(\bm{x},\bm{x}^{(i)}\right)\tilde{\rho}_{T}(\bm{x}^{(i)}),\quad\text{with}\leavevmode\nobreak\ \kappa(\bm{x},\bm{x}^{(i)})=\sum_{\bm{k}\in\mathbb{Z}^{d},\|\bm{k}\|_{2}\leq\Lambda}\cos\left(\pi\bm{k}\cdot\left(\bm{x}-\bm{x}^{(i)}\right)\right)\in\mathbb{R},$ (138)
@@
Since $P_{i}$, $P_{j}$, and $P_{k}$ are Pauli strings, any two of them must commute or anti-commute to each other. By exploiting this property, we next separately analyze the value of $[\operatorname{\mathfrak{R}_{P}}_{\bm{a}_{j}}(\bm{x}_{j})]_{ik}$ for $\forall i,k\in[4^{d}]$.
Case I: $[P_{i},P_{j}]=0$ or $[P_{k},P_{j}]=0$. In this case, Eq. (140) can be simplified to
$[\operatorname{\mathfrak{R}_{P}}_{\bm{a}_{j}}(\bm{x}_{j})]_{ik}=\operatorname{Tr}(P_{i}P_{k})=\delta_{ij}.$ (141)
In other words, for the indices $i,k\in[4^{d}]$ whose corresponding Pauli strings commute with $P_{j}$, the corresponding PTM entry is a constant, which is zero if $i\neq k$ and $1$ if $i=k$. According to the formula in Eq. (139), the relevant entries, whose indices satisfy $[P_{i},P_{j}]=0$ or $[P_{k},P_{j}]=0$, belong to $D_{0}^{\prime}$.
@@
$[\operatorname{\mathfrak{R}_{P}}_{\bm{a}_{j}}(\bm{x}_{j})]_{ik}=\operatorname{Tr}(\operatorname{RP}_{j}(2x)P_{i}P_{k})=\cos(x)\operatorname{Tr}(P_{i}P_{k})+\imath\sin(x)\operatorname{Tr}(P_{j}P_{i}P_{k})=\cos(x)\delta_{ij}+\imath\sin(x)\operatorname{Tr}(P_{j}P_{i}P_{k}),$ (142)

### v2

@@ A.4 Literature review
$\rho_{\Lambda}=\sum_{\bm{\omega}\in\mathfrak{C}(\Lambda)}\Phi_{\bm{\omega}}(\bm{x})\rho_{\bm{\omega}}\penalty 10000\ \textnormal{and}\penalty 10000\ \rho=\sum_{\bm{\omega}\in\mathfrak{C}(d)}\Phi_{\bm{\omega}}(\bm{x})\rho_{\bm{\omega}},$ (86)
@@
$\Psi_{\bm{\omega}_{i}}(\bm{x}_{i})\equiv\nabla_{\bm{x}_{i}}\Phi_{\bm{\omega}_{i}}(\bm{x}_{i})=\begin{cases}0\penalty 10000\ &\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=0\\ -\sin(\bm{x}_{i})&\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=1\\ \cos(\bm{x}_{i})&\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=-1\end{cases}.$ (90)
@@
where Eqs. (91h)-(91h) are obtained by taking the expectation value on the gradient term $\Psi_{\bm{\omega}_{i}}$ for $\forall i\in[d]$, the quantity $\alpha_{\pm 1,\bm{\omega}_{2:d}}$ refers to $\operatorname{Tr}(\rho_{\pm 1,\bm{\omega}_{2:d}}O)$ with the frequency value at the $1$-st position being $\pm 1$ (the same rule applies to other terms such as $\alpha_{\bm{\omega}_{1:i},i+1,\bm{\omega}_{i+2:d}}$), and Eq. (91j) is obtained by taking integral over each term in the bracket and the orthogonality of basis functions in $\Phi_{\bm{\omega}}(\bm{x})$ and $\Phi_{\bm{\omega}^{\prime}}(\bm{x})$ as shown in Eq. (88).
@@
Note that depending on the number of non-zero entries of the frequency (i.e., $\|\bm{\omega}\|_{0}$), the left and right-hand sides in Eq. (93) can be decomposed into $d-\Lambda$ parts and d1d-1 parts, respectively. Mathematically, the left hand side of Eq. (93) yields
@@
$\bigstar=$ $\Biggl[\bigg(\sum_{\bm{\omega}_{2:d},\|\bm{\omega}_{2:d}\|_{0}=0}2^{-\|[1,\bm{\omega}_{2:d}]\|_{0}}|\alpha_{1,\bm{\omega}_{2:d}}|^{2}+\cdots+\sum_{\bm{\omega}_{2:d},\|\bm{\omega}_{2:d}\|=d-1}2^{-\|[1,\bm{\omega}_{2:d}]\|_{0}}|\alpha_{1,\bm{\omega}_{2:d}}|^{2}\bigg)$ (97)
$+\cdots$
@@
$\Lambda^{\prime}\sum_{\bm{\omega},\|\bm{\omega}\|_{0}=\Lambda^{\prime}}2^{-\|\bm{\omega}\|_{0}}|\alpha_{\bm{\omega}}|^{2},\quad\text{with}\penalty 10000\ \Lambda<\Lambda^{\prime}<d.$ (102)
@@
Supplementary Fig. 2: A visual interpretation about the appearance times of $\tilde{\bm{\omega}}$ in the right-hand side of Eq. (101).
Combining with the above two equations, we know that the total number of frequencies in Eq. (102) and the right-hand side of Eq. (101) is the same. Accordingly, their equivalence can be reached if we can show that each frequency $\bm{\omega}\in\{0,\pm 1\}^{d}$ with $\|\bm{\omega}\|_{0}=\Lambda^{\prime}$ appears $\Lambda^{\prime}$ times in the right-hand side of Eq. (101). This is indeed the case, as the visual interpretation is shown in Fig. 2. Denote that the specified frequency as $\tilde{\bm{\omega}}$ with $\|\tilde{\bm{\omega}}\|_{0}=\Lambda^{\prime}$. In other words, there are $d-\Lambda^{\prime}$ entries whose value is 0. This property allows us to determine how many times $\tilde{\bm{\omega}}$ appears on the right side of Eq. (101). In particular, as shown in the left panel of Fig. 2, the right-hand side of Eq. (101) can be divided into $d$ groups, depending on the location of $\pm 1$. Note that among these $d$ groups, every entry with the value 0 in $\tilde{\bm{\omega}}$ precludes one group, and only $d-(d-\Lambda^{\prime})=\Lambda^{\prime}$ feasible groups are preserved, as shown in the right panel of Fig. 2. Moreover, for each feasible group, $\tilde{\bm{\omega}}$ can only appear once, because of the orthogonality of different frequencies. Consequently, we obtain that each frequency $\tilde{\bm{\omega}}\in\{0,\pm 1\}^{d}$ with $\|\tilde{\bm{\omega}}\|_{0}=\Lambda^{\prime}$ appears $\Lambda^{\prime}$ times, indicating that the right-hand side in Eq. (101) equals to Eq. (102).
@@
[…]
@@
where the first inequality uses the Cauchy–Schwarz inequality and $|\Phi_{\bm{\omega}}(\bm{x})|\leq 1$, the second inequality adopts the von Neumann’s trace inequality with Hölder’s inequality, and the last inequality exploits $\|\tilde{\rho}_{1}(\bm{x})\|_{1}\leq 3^{K}$ [83, Eq. (F37)] and the condition $\|O\|_{\infty}\leq B$.
@@
Denote the set of truncated frequencies as $\mathfrak{C}(\Lambda)=\{\bm{\omega}|\bm{\omega}\in\{0,\pm 1\}^{d},\penalty 10000\ s.t.\penalty 10000\ \|\bm{\omega}\|_{0}\leq\Lambda\}$. This leads to
@@
Here we analyze the computational cost of the proposed ML model. For clarity, here we separately analyze the computational time of our proposal required in the training and inference.
@@ H.1 Implementation of the classical prediction model
Supplementary Fig. 3: Purely classical learning model with the full expansion. The pipeline is similar to the one presented in Fig. 1 of the main text. The only difference is the way of collecting training data, where the circuit layout is known by the learner and the expectation value $\operatorname{Tr}(\rho(bx)O)$ should be efficiently calculated by the employed classical simulators.
The implementation of the protocol introduced in Corollary 1 is visualized in Fig. 3. The only difference with the one introduced in the main text is the way of collecting training data. Specifically, when the tunable quantum circuit architecture formed by $\operatorname{RZ}+\operatorname{\text{CI}}$ is known to the learner, the classical simulators (e.g., tensor network simulators, LOWESA, and near Clifford circuits simulators) are employed to collect the training data
@@
$h_{\mathsf{c}}(\bm{x},O)=\frac{1}{n}\sum_{i=1}^{n}\kappa\left(\bm{x},\bm{x}^{(i)}\right)\operatorname{Tr}(\rho(\bm{x}^{(i)})O)\penalty 10000\ \text{with}\penalty 10000\ \kappa(\bm{x},\bm{x}^{(i)})=\sum_{\bm{\omega},\|\bm{\omega}\|_{0}\leq d}2^{\|\bm{\omega}\|_{0}}\Phi_{\bm{\omega}}(\bm{x})\Phi_{\bm{\omega}}(\bm{x}^{(i)})\in\mathbb{R}.$ (126)
@@ H.2 Proof of Corollary 1
SM I Learning bounded-gate quantum circuit with $\operatorname{\text{CI}}$ gates and parameterized multi-qubit gates
In the main text, we primarily elaborate on the application of the proposed ML model in predicting bounded-gate quantum circuits consisting of $\operatorname{RZ}$ and $\operatorname{\text{CI}}$ gates. However, our proposal is adaptable and can be readily extended to predict bounded-gate quantum circuits composed of alternative basis gate sets. To exemplify this flexibility, in this section, we illustrate how our proposal and the associated theoretical findings (Theorem 2) can be effectively expanded to encompass a wider context, specifically bounded-gate circuits incorporating CI gates alongside parameterized multi-qubit gates generated by arbitrary Pauli strings.
@@
[…]
where the trigonometric monomial basis is $\Phi_{\bm{\omega}}(\bm{x})=\prod_{i=1}^{d}\begin{cases}1\penalty 10000\ &\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=0\\ \cos(\bm{x}_{i})&\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=1\\ \sin(\bm{x}_{i})&\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=-1\end{cases}$, then we can form the state prediction model in Eq. (76), i.e.,
$\hat{\sigma}_{n}(\bm{x})=\frac{1}{n}\sum_{i=1}^{n}\kappa\left(\bm{x},\bm{x}^{(i)}\right)\tilde{\rho}_{T}(\bm{x}^{(i)}),\quad\text{with}\penalty 10000\ \kappa(\bm{x},\bm{x}^{(i)})=\sum_{\bm{k}\in\mathbb{Z}^{d},\|\bm{k}\|_{2}\leq\Lambda}\cos\left(\pi\bm{k}\cdot\left(\bm{x}-\bm{x}^{(i)}\right)\right)\in\mathbb{R},$ (139)
@@
Since $P_{i}$, $P_{j}$, and $P_{k}$ are Pauli strings, any two of them must commute or anti-commute to each other. By exploiting this property, we next separately analyze the value of $\operatorname{\mathfrak{R}_{P}}_{\bm{a}_{j}}(\bm{x}_{j})]_{ik}$ for $\forall i,k\in[4^{d}]$.
Case I: $P_{i},P_{j}]=0$ or $P_{k},P_{j}]=0$. In this case, Eq. (141) can be simplified to
$\operatorname{\mathfrak{R}_{P}}_{\bm{a}_{j}}(\bm{x}_{j})]_{ik}=\operatorname{Tr}(P_{i}P_{k})=\delta_{ij}.$ (142)
In other words, for the indices $i,k\in[4^{d}]$ whose corresponding Pauli strings commute with $P_{j}$, the corresponding PTM entry is a constant, which is zero if $i\neq k$ and $1$ if $i=k$. According to the formula in Eq. (140), the relevant entries, whose indices satisfy $P_{i},P_{j}]=0$ or $P_{k},P_{j}]=0$, belong to $D_{0}^{\prime}$.
@@
$\operatorname{\mathfrak{R}_{P}}_{\bm{a}_{j}}(\bm{x}_{j})]_{ik}=\operatorname{Tr}(\operatorname{RP}_{j}(2x)P_{i}P_{k})=\cos(x)\operatorname{Tr}(P_{i}P_{k})+\imath\sin(x)\operatorname{Tr}(P_{j}P_{i}P_{k})=\cos(x)\delta_{ij}+\imath\sin(x)\operatorname{Tr}(P_{j}P_{i}P_{k}),$ (143)